### v1

# RDDM: Practicing RAW Domain Diffusion Model for Real-world Image Restoration

Abstract
We present the RAW domain diffusion model (RDDM), an end-to-end diffusion model that restores photo-realistic images directly from the sensor RAW data. While recent sRGB-domain diffusion methods achieve impressive results, they are caught in a dilemma between high fidelity and realistic generation. As these models process lossy sRGB inputs and neglect the accessibility of the sensor RAW images in many scenarios, e.g., in image and video capturing in edge devices, resulting in sub-optimal performance. RDDM bypasses this limitation by directly restoring images in the RAW domain, replacing the conventional two-stage image signal processing (ISP)IR pipeline. However, a simple adaptation of pre-trained diffusion models to the RAW domain confronts the out-of-distribution (OOD) issues. To this end, we propose: (1) a RAW-domain VAE (RVAE) learning optimal latent representations, (2) a differentiable Post Tone Processing (PTP) module enabling joint RAW and sRGB space optimization. To compensate for the deficiency in the dataset, we develop a scalable degradation pipeline synthesizing RAW LQ-HQ pairs from existing sRGB datasets for large-scale training. Furthermore, we devise a configurable multi-bayer (CMB) LoRA module handling diverse RAW patterns such as RGGB, BGGR, etc. Extensive experiments demonstrate RDDM’s superiority over state-of-the-art sRGB diffusion methods, yielding higher fidelity results with fewer artifacts.
1 Introduction
Real-world Image Restoration (Real-IR) aims to restore high-quality (HQ) images from low-quality (LQ) images containing complex degradations, e.g. noise, image compression and blur (Fan et al., 2020; Jinjin et al., 2020; Zhang et al., 2022; 2019; 2018b; 2017). Existing GAN-based (Ledig et al., 2017; Wang et al., 2018) methods employ a generator and a discriminator for adversarial training. However, GAN-based methods suffer from pattern collapse, incurring unsatisfactory results (Wang et al., 2021; Zhang et al., 2021; Liang et al., 2022a; Chen et al., 2022; Liang et al., 2022b; Xie et al., 2023). Benefiting from the powerful generative priors granted by text-to-image (T2I) models, SUPIR (Yu et al., 2024) and its counterparts (Zhang et al., 2023; Lin et al., 2024; Wang et al., 2024; Wu et al., 2024b; Yu et al., 2024; Sun et al., 2024; Menon et al., 2020; Karras et al., 2019) integrate T2I models into Real-IR and have attained remarkable performance. Nevertheless, the dilemma between image fidelity and realistic generation remains a pivotal challenge that these methods must confront.
Notably, prevailing Real-IR models process and enhance images in the lossy sRGB domain. This may lead to sub-optimal results in many scenarios where the sensor RAW images are accessible, e.g., in image and video capturing in edge devices, as the rich meta-info contained in the sensor RAW images is not effectively squeezed. This drives us to integrate the powerful pre-trained T2I models with sensor RAW data.
However, a naive combination of prior models trained in the sRGB domain with the sensor RAW images confronts substantial challenges. Firstly, sRGB images and sensor RAW images differ significantly in terms of luminance, mosaic patterns of RAW images and noise distribution, as shown in Fig. 1, so that publicly available VAE pre-trained in the sRGB domain cannot effectively encode and decode sensor RAW images. Secondly, a comprehensive RAW domain Real-IR dataset that could serve as a rigorous benchmark for training and evaluating IR models in the RAW domain is still absent.
To address these issues, in this paper we propose the RAW domain diffusion model (RDDM). As a remedy to the first challenge, we devise a RAW domain VAE (RVAE) that accepts a sensor RAW as the input and outputs a clean image in the linear domain. We employ a divide-and-conquer strategy to train the RVAE. Initially, we fine-tune a pre-trained VAE using a linear HQ dataset to bridge the gap between the sRGB and linear domains. Subsequently, we adopt the LoRA fine-tuning approach to jointly train the encoder of the RVAE and Real-IR, thereby adapting it to the mosaic pattern of real sensor RAW data. To further mitigate out-of-distribution (OOD) issues, we design a differentiable post tone processing (PTP) module that enables joint RAW and sRGB space optimization, thereby endowing the model with improved fidelity. Additionally, a configurable multi-Bayer (CMB) LoRA module is designed to adapt our model to different Bayer patterns of RAW. To compensate for the deficiency in the dataset, we propose a realistic RAW domain degradation pipeline that amalgamates the degradation strategies of ESRGAN (Wang et al., 2021) and UPI (Brooks et al., 2019), allowing us to synthesize abundant RAW–linear image pairs from publicly accessible sRGB datasets.
In summary, our main contributions are as follows:
We propose RDDM, the first practical application of the raw domain diffusion model, establishing a novel paradigm for RAW image restoration.
We propose RVAE capable of encoding mosaiced RAW images and subsequently decoding the latent representations into linear HQ images in order to address the OOD issues.
Additionally, we design a RAW domain Real-IR data synthesis pipeline and construct a RAW Real-IR benchmark.
Extensive experiments verify that RDDM demonstrates superior image fidelity and comparable generation capability to the state-of-the-art methods.
2 Related Work
Real-world Image Restoration. Real-IR is becoming a trending field of research since the advent of ESRGAN (Ledig et al., 2017). Early studies attempted various ways to combine generative adversarial networks (GANs) (Goodfellow et al., 2014; Karras et al., 2017; 2019; Radford et al., 2015; Mirza & Osindero, 2014) with perceptual losses (Ding et al., 2020; Johnson et al., 2016; Zhang et al., 2018a) for training networks to predict images that follow the natural image distribution (Ledig et al., 2017; Wang et al., 2018; 2021; Zhang et al., 2021; Liang et al., 2022a; Chen et al., 2022; Liang et al., 2022b; Xie et al., 2023). However, since adversarial training of GANs can be unstable, their discriminators are deficient in determining the quality of the diverse natural image contents, giving rise to unnatural visual artifacts. As an alternative to GAN-based methods, diffusion-based models (Podell et al., 2023; Rombach et al., 2022) is becoming increasingly popular in Real-IR tasks (Kawar et al., 2022; Li et al., 2022; Luo et al., 2023a; b; Özdenizci & Legenstein, 2023; Saharia et al., 2022) to generate realistic images with substantial texture, leveraging pre-trained Stable Diffusion (SD) models as priors (Zhang et al., 2023; Lin et al., 2024; Wang et al., 2024; Wu et al., 2024b; Yu et al., 2024; Sun et al., 2024; Menon et al., 2020; Karras et al., 2019) whereas they employ different condition injection strategies and feature extraction. Nevertheless, all existing Real-IR methods restore images in the sRGB domain in which rich information in the RAW domain might be lost after ISP. However, directly adapting sRGB Real-IR methods to the RAW domain encounters severe domain mismatch and results in poor performance.
Sensor RAW Images. Sensor RAW refers to the unprocessed, original data collected directly from specific camera sensors, holding a wealth of physical information. Sensor RAW contains noise and the Bayer Color Filter Array (CFA) patterns (Snyder et al., 1995; Beenakker & Patra, 1999; Gotoh & Okutomi, 2004; Maschal Jr et al., 2010). The demosaicing process removes the Bayer CFA pattern and produces linear domain images from RAW, and the Post Tone Processing (PTP) module produces natural sRGB image from linear domain images. In short, sensor RAW images differ from sRGB ones in terms of sensor-captured information (e.g. 12-16 bit photon-electric signals), noise distribution, color space, luminance, dynamic range and image format (e.g. mosaic patterns), making it difficult for existing sRGB Real-IR methods to adapt to the RAW domain.
Image Signal Processing. An ISP pipeline reconstructs a visually appealing sRGB image from RAW sensor data. Traditionally, an ISP pipeline is formulated as a series of hand-crafted modules executed sequentially (Sundararajan, 2017), including some representative steps such as demosaicing (DM), denoising (DN), automated-white-balance (AWB), color correction matrix (CCM), gamma compression (GC), and tone mapping (TM). Among these steps, DM (Li et al., 2008; Alleysson et al., 2005; Hirakawa & Parks, 2005; Kimmel, 1999) and DN modules (Brooks et al., 2019; Cao et al., 2024; Li et al., 2024) are ill-posed and can cause over-smoothen images (Qian et al., 2019) and AWB, CCM and TM modules can cause information compression due to their inherent data clipping operations. In addition, different digital imaging device producers adopt different ISP pipelines which usually remain as black boxes, making it difficult to obtain information about the specific steps inside.
3 Methodology
3.1 Preliminaries
Problem Modeling. Real-world image restoration in the sRGB domain aims to train a neural network , parameterized by , estimating HQ sRGB image given LQ sRGB image . In RAW domain, we train a neural network to transform LQ sensor RAW to HQ linear domain image . The training task can be modeled as the following optimization problem:
where S is the dataset consisting of () pairs, and is the loss function, respectively.
Image Signal Proccessing. Given a sensor RAW with single channel as input, DN and DM module produces linear domain image with three channels, which can be formulated as:
where is the inverse function, transfers the linear domain image to the RAW domain image . Post Tone Processing (PTP) module including AWB, CCM, GC and TM, converts linear domain images to sRGB images , which is defined as:
where strips the color information from an sRGB-encoded image and converts the resulting grayscale representation into the linear domain. A detailed introduction of ISP and Inverse ISP, as well as their mathematical derivations referenced in the Appendix.
3.2 RAW Domain Diffusion Model
Framework Overview. Our generator is composed of an RVAE encoder , a diffusion network , and an RVAE decoder . extracts the latent features of the sensor RAW image. jointly optimizes DN, DM, and image restoration in the latent space to obtain the latent features. then decodes the latent features to produce the linear domain image, which is subsequently mapped to the sRGB domain by the PTP module. We incorporate trainable LoRA layers (Hu et al., 2022) into the pre-trained and . To address the issue of extracting prompts from sensor RAW images, the feed-forward ISP firstly processes the sensor RAW image to obtain an sRGB image, from which the DAPE (Zheng et al., 2024) prompt extractor extracts the textual information to activate priors in the model training. Fig. 3 presents an overview of the framework and the interplay between the various modules.
RAW Domain VAE. VAE plays a pivotal role in the quality of generated images. However, existing VAE in the sRGB domain are not capable of effectively encoding RAW images and decoding linear domain images. Therefore, we train a RAW domain VAE that encodes RAW images and subsequently decodes the latent representation into a linear domain image. (Rombach et al., 2022) employs a scaling factor to normalize the latent space distributions of different VAEs to the standard Gaussian distribution, which is beneficial for diffusion network optimization. To obtain the statistically accurate scaling factor, we calculate the parameter for training samples in the linear domain according to the following formula:
where denotes the latent space of the training samples encoded by . and present the mean and variance of the data distribution. The rescaled latent has unit standard deviation, i.e., .
The training strategy of RVAE is illustrated in Fig. 3 (a). We train the encoder and decoder on the linear domain dataset, such that the linear domain input is encoded by RVAE encoder to obtain the latent feature and the RVAE decoder decodes into the target linear domain image . Furthermore, we introduce a differentiable PTP module that simultaneously supervises training in both the sRGB and RAW domains. Similar to LDM (Rombach et al., 2022), we use loss, LPIPS loss, and GAN loss to train the VAE encoder and decoder to generate realistic details of a linear image:
where and is calculated in both the RAW and sRGB domains. and represents the gradient of the last layer in the decoder.
Training Framework. In order to accommodate RAW images captured with arbitrary Bayer patterns, we propose a configurable multi-Bayer (CMB) LoRA module that augments the RVAE encoder and the pre-trained diffusion network with independent sets of LoRA, and we assign a distinct LoRA group to each Bayer pattern. During training, the RVAE decoder is frozen and only the CMB modules are optimized. Consequently, the encoder learns to encode RAW images and diffusion network jointly performing DN, DM, and detail enhancement. We use VSD loss, LPIPS loss, and MSE loss to train our model in the RAW domain and the sRGB domain:
where , are weighting scalars. . . We transform both the model predictions and the ground-truth into the sRGB domain via the proposed PTP module.
Feed-forward ISP. To supervise the training of RDDM in sRGB domain and extract prompt information from RAW images, we devise a feed-forward ISP that can transform RAW images into the sRGB domain, as illustrated in Fig. 4 (b). The required metadata, including Bayer pattern, AWB gain, and CCM, are obtained from Inverse ISP and subsequently employed as parameters of the feed-forward ISP. We train a lightweight neural network DDNet to replace conventional DN and DM modules by minimizing: , where is the joint DN and DM network, is the sensor RAW and is the HQ linear image.
RAW prompt extractor. To address the issue that existing text extractors fail to accurately extract text information from sensor RAW images with significant noise, we propose the RAW prompt extractor (RPE). Initially, we train a lightweight denoising and demosaicing network to transform noisy sensor RAW images into clean linear domain ones. Subsequently, the PTP module converts these linear domain images into sRGB images. Finally, we employ the DAPE to accurately extract the prompt information. can significantly reduce the impact of noise on text extractors, thereby enhancing the robustness of the prompt extractor model.
3.3 Image Synthesis Pipeline
Synthetic Image Degradation Pipeline. Despite the abundance of existing datasets for Real-IR, such as LSDIR (Li et al., 2023), FFHQ (Karras et al., 2019), and DIV2K (Agustsson & Timofte, 2017), these datasets are all in the sRGB domain. To the best of our knowledge, there is currently no dataset for Real-IR in the RAW domain. Therefore, to provide a solid training foundation for Real-IR in the RAW domain, we synthesize a RAW domain Real-IR dataset by degrading publicly available sRGB Real-IR datasets through our synthetic image degradation pipeline, as shown in Fig. 4 (a). In particular, we first degrade sRGB HQ images to sRGB LQ images via detail degradation method, following Real-ESRGAN (Wang et al., 2021) despite excluding the degradation process of random noise, since the noise in the RAW domain is intrinsic in the sensor’s physical features. Subsequently, for the synthesis of the training dataset for RDDM, the sRGB LQ images are processed through the inverse post tone processing module (IPTP) and mosaic noise synthesizer module (MNS) to obtain degraed RAW images , while the sRGB HQ images are transformed into linear domain GT through inverse PTP module. For the synthesis of the training dataset for DDNet, linear HQ images are processed through MNS module to produce detailed RAW images .
4 Experiments
| Dataset | Method | PSNR | SSIM | LPIPS | DISTS | FID | NIQE | MUSIQ | CLIPIQA |
| DIV2K-Val | JDnDmSR | 23.4565 | 0.6192 | 0.5347 | 0.2655 | 45.3706 | 7.0895 | 32.1252 | 0.1978 |
| SwinIR | 22.7983 | 0.6294 | 0.5345 | 0.2780 | 44.9270 | 7.1012 | 32.9053 | 0.2520 | |
| ISP+StableSR-s200 | 23.6034 | 0.6133 | 0.4095 | 0.2092 | 35.6300 | 4.7840 | 43.8325 | 0.4284 | |
| ISP+DiffBIR-s50 | 22.4903 | 0.5284 | 0.4519 | 0.2176 | 42.0167 | 4.6040 | 52.9640 | 0.6503 | |
| ISP+PASD-s20 | 23.3860 | 0.6150 | 0.3029 | 0.1385 | 23.5801 | 3.4392 | 64.3181 | 0.6197 | |
| ISP+SeeSR-s50 | 23.2836 | 0.6059 | 0.2880 | 0.1363 | 25.4424 | 3.5605 | 65.6650 | 0.6976 | |
| ISP+SUPIR-s50 | 22.4837 | 0.5935 | 0.3265 | 0.1462 | 27.4418 | 3.5376 | 62.7078 | 0.5570 | |
| ISP+OSEDiff-s1 | 22.5277 | 0.6069 | 0.2836 | 0.1351 | 38.0461 | 3.6427 | 66.2024 | 0.6818 | |
| Ours-s1 | 23.7416 | 0.6296 | 0.2540 | 0.1197 | 23.8028 | 3.3627 | 65.4202 | 0.6737 | |
| DRealSR | JDnDmSR | 27.6972 | 0.7995 | 0.3610 | 0.2210 | 31.1697 | 7.9294 | 30.4728 | 0.2373 |
| SwinIR | 27.0657 | 0.8161 | 0.3714 | 0.2305 | 30.5639 | 7.5234 | 30.2268 | 0.2972 | |
| ISP+StableSR-s200 | 27.1173 | 0.7613 | 0.3387 | 0.1978 | 25.8442 | 4.5959 | 49.2604 | 0.5991 | |
| ISP+DiffBIR-s50 | 28.2670 | 0.7606 | 0.4142 | 0.2702 | 25.9530 | 6.3725 | 38.1396 | 0.5284 | |
| ISP+PASD-s20 | 28.3377 | 0.7845 | 0.2870 | 0.1670 | 16.1714 | 4.6875 | 53.1539 | 0.5872 | |
| ISP+SeeSR-s50 | 27.6513 | 0.7765 | 0.2972 | 0.1816 | 19.2938 | 4.2053 | 56.0800 | 0.6681 | |
| ISP+SUPIR-s50 | 26.9559 | 0.7359 | 0.3262 | 0.1799 | 26.1866 | 5.0892 | 48.5114 | 0.4839 | |
| ISP+OSEDiff-s1 | 25.1101 | 0.7315 | 0.3396 | 0.1900 | 32.4002 | 4.7336 | 57.3375 | 0.7376 | |
| Ours-s1 | 28.3495 | 0.7892 | 0.2719 | 0.1649 | 17.4825 | 4.6852 | 57.0696 | 0.7035 | |
| RealSR | JDnDmSR | 25.6346 | 0.7532 | 0.3649 | 0.2119 | 66.5709 | 7.4103 | 38.6661 | 0.2062 |
| SwinIR | 25.4564 | 0.7477 | 0.3818 | 0.2283 | 67.1467 | 6.9218 | 38.5181 | 0.2637 | |
| ISP+StableSR-s200 | 23.3339 | 0.6600 | 0.3505 | 0.1949 | 60.9322 | 3.9343 | 64.1478 | 0.6393 | |
| ISP+DiffBIR-s50 | 25.3643 | 0.6761 | 0.4086 | 0.2478 | 56.7401 | 5.6140 | 49.4878 | 0.5581 | |
| ISP+PASD-s20 | 24.8545 | 0.6886 | 0.3055 | 0.1720 | 40.8756 | 4.1290 | 63.5759 | 0.6223 | |
| ISP+SeeSR-s50 | 24.8332 | 0.6957 | 0.2872 | 0.1807 | 36.0702 | 4.2017 | 66.3191 | 0.6977 | |
| ISP+SUPIR-s50 | 23.9782 | 0.6505 | 0.3412 | 0.1937 | 51.7890 | 4.9086 | 59.3107 | 0.4814 | |
| ISP+OSEDiff-s1 | 23.8067 | 0.6872 | 0.2988 | 0.1768 | 52.0761 | 4.2011 | 65.5805 | 0.6793 | |
| Ours-s1 | 25.1264 | 0.7092 | 0.2546 | 0.1589 | 36.8671 | 4.1286 | 65.8881 | 0.6723 |
4.1 Experimental Settings
Training and Testing Datasets. We adopt the OSEDiff (Wu et al., 2024a) setup and train RDDM using the LSDIR (Li et al., 2023) dataset and the first 10K face images from FFHQ (Karras et al., 2019). We use the degradation pipeline discussed in the Image Synthesis Pipeline section to synthesize LQ and HQ pairs in the RAW domain. For testing, we construct our benchmark by degrading the HR images from the DIV2K-Val, consisting of 100 images, RealSR containing 100 images, and DRealSR containing 93 images, using our proposed data synthesis pipeline.
| Method | Params(M) | FLOPs(G) |
| JDnDmSR | 78.2 | 54 |
| SwinIR | 11.605 | 53 |
| ISP+StableSR-s200 | 1413 | 830 |
| ISP+DiffBIR-s50 | 1673 | 1670 |
| ISP+PASD-s20 | 1432 | 1590 |
| ISP+SeeSR-s50 | 1622 | 1230 |
| ISP+SUPIR-s50 | 4805 | 4100 |
| ISP+OSEDiff-s1 | 1298 | 250 |
| Ours-s1 | 1294 | 250 |
Compared Methods. We compare RDDM against best best-performing traditional one-stage method and two-stage ISPIR models methods, as shown in Table 1. For one-stage methods, we choose JDnDmSR (Xing & Egiazarian, ) and SwinIR (Liang et al., 2021) as our baseline. For two-stage ISPIR methods, we choose PIPNet (A Sharif et al., 2021) as the DN and DM module for ISP and diffusion-based IR methods as our Real-IR baselines.
Evaluation Metrics. For a thorough assessment of the different methods, we utilize a variety of full-reference and non-reference evaluation metrics to test each method’s image fidelity and generation quality. PSNR and SSIM (Wang et al., 2004) (calculated on 3 channels) measure image fidelity, whereas LPIPS (Zhang et al., 2018a) and DISTS (Ding et al., 2020) measure perceptual qualities based on reference images. FID (Heusel et al., 2017) assesses the distributional distance between the GT and the restored images. NIQE (Mittal et al., 2012), MUSIQ (Ke et al., 2021), and CLIPIQA (Wang et al., 2023) are non-reference image generation quality measurements.
Implementation Details. We train RDDM with the AdamW optimizer at a learning rate of . The entire training process spans 150000 steps with a batch size of 16. The rank of LoRA in the RVAE Encoder and the diffusion network is set to 4. We employ DAPE as the sRGB domain text prompts extractor.
4.2 Comparisons with State-of-the-Arts
Quantitative Comparisons. The quantitative comparisons among the competing methods on the three test datasets are presented in Table 1. Compared against other approaches, RDDM ranks top 3 in terms of all metrics, including full-reference fidelity metrics like PSNR, SSIM and perceptual quality metrics such as LPIPS and DISTS, on DIV2K-Val, DRealSR and RealSR datasets, except for only PSNR on RealSR. While JDnDmSR and SwinIR achieve slightly higher PSNR and SSIM scores than RDDM on the RealSR dataset, they exhibit markedly inferior performance on other metrics, particularly on no-reference metrics like NIQE, MUSIQ, and CLIPIQA, and distribution alignment metric FID. This suggests that their generative capabilities are substantially weaker than RDDM. In addition, RDDM achieves comparable generative performance with diffusion-based methods, while considerably outperforming them from the aspects of image fidelity metrics such as PSNR and SSIM. Overall, RDDM achieves superior image fidelity and comparable generation capability to the state-of-the-art methods. Note that Table 4.1 further shows the Params and FLOPs of the competing methods. While presenting adequate performace, RDDM has the least Params and FLOPs among diffusion-based models.
Qualitative Comparison. Fig. 5 presents the visual comparisons of RDDM on RealSR, DIV2K-val and DRealSR, along with traditional one-stage and two-stage methods. In our first example, the wall textures generated by JDnDmSR, SwinIR, and DiffBIR are notably blurry, with JDnDmSR exhibiting a significant color deviation. PASD, SeeSR, StableSR, and OSEDiff produce more textures but still fail to capture finer details. SUPIR generates clearer results, yet the textures appear highly unnatural with numerous artifacts. RDDM effectively leverages the detailed information from sensor RAW data to produce more realistic wall textures with higher clarity. The second example draws the same conclusion. It indicates that fully utilizing the abundant yet often lost information in sensor RAW can effectively address the artifact issue prevalent in existing diffusion-based methods. The third and fourth comparisons further manifests that RDDM is capable of generating clear textures with finer details, higher fidelity and less noise. More visualization comparison results can be found in the Appendix. To further investigate the user preferences about these results, we conduct a user study comparing our method on RealSR, DrealSR, and DIV2K-val test datasets, with 5 participants involved. For each set of comparison images, users select their preferred result. As shown in Fig. 6, the results demonstrate that our method significantly outperforms state-of-the-art methods in terms of perceptual quality.
| PSNR | SSIM | LPIPS | DISTS | FID | NIQE | MUSIQ | CLIPIQA | |
| sRGB VAE | 22.2130 | 0.6200 | 0.3654 | 0.2035 | 58.9168 | 4.4919 | 63.6258 | 0.6912 |
| RVAE | 25.1264 | 0.7092 | 0.2546 | 0.1589 | 36.8671 | 4.1286 | 65.8881 | 0.6723 |
4.3 Ablation Study
| RAW Loss | sRGB Loss | PSNR | SSIM | LPIPS | DISTS | FID | NIQE | MUSIQ | CLIPIQA |
| ✓ | 22.9510 | 0.6096 | 0.2794 | 0.1337 | 36.3172 | 3.4637 | 66.1391 | 0.7298 | |
| ✓ | 23.3129 | 0.6237 | 0.2635 | 0.1216 | 29.5131 | 3.3352 | 64.7748 | 0.6637 | |
| ✓ | ✓ | 23.7416 | 0.6296 | 0.2540 | 0.1197 | 23.8028 | 3.3627 | 65.4202 | 0.6737 |
| Setting | PSNR | SSIM | LPIPS | FID |
| 1 | 25.2625 | 0.8017 | 0.1719 | 41.8118 |
| 2 | 27.1487 | 0.7035 | 0.3068 | 48.4675 |
| 3 | 27.6892 | 0.7176 | 0.2787 | 38.2007 |
| 4 | 32.5424 | 0.9082 | 0.0533 | 11.6156 |
The importance of RVAE. To illustrate the importance of the proposed RVAE, we substitute it with a pre-trained sRGB VAE. The quantitative comparison on RealSR testset is shown in Table 3. The RDDM employing RVAE outperforms its counterpart utilizing a pre-trained sRGB VAE in both fidelity metrics and perceptual quality metrics with the exception of CLIPIQA. Fig. 7 illustrates that RVAE can significantly mitigate the color deviation issue in RDDM. This is attributed to the fact that the pre-trained sRGB VAE is incapable of adapting to sensor RAW images. In contrast, our RVAE is capable of effectively encoding the sensor RAW data and decoding it into the linear domain images.
| PSNR | SSIM | LPIPS | DISTS | FID | NIQE | MUSIQ | CLIPIQA | |
| sRGBPE | 24.7811 | 0.7201 | 0.2495 | 0.1601 | 38.1092 | 4.2538 | 65.4506 | 0.6579 |
| ISPPE | 24.7816 | 0.7201 | 0.2495 | 0.1600 | 38.0969 | 4.2539 | 65.4526 | 0.6579 |
| RPE | 25.1264 | 0.7092 | 0.2546 | 0.1589 | 36.8671 | 4.1286 | 65.8881 | 0.6723 |
Setting of RAW Domain VAE. We systematically evaluate four transfer strategies: (1) We directly use a pre-trained sRGB domain VAE. The model fails to reconstruct the regular Bayer mosaic present in the sensor RAW data and is unable to produce a faithful linear domain image, as depicted in Fig. 8(a). (2) We construct the end-to-end training of the VAE with sensor RAW and linear image pairs. We observe that the model tends to produce images with significant mosaic textures, as illustrated in Fig. 8(b). (3) We first train the encoder and decoder with linear domain images, and then freeze the decoder while training the LoRA layers applied to the encoder with sensor RAW and linear image pairs. However, the result is still dominated by noticeable mosaic textures, as depicted in Fig. 8(c). (4) Following the same first training stage as (3), we freeze the decoder and conduct joint training by incorporating LoRA layers into both the encoder and the diffusion network. As shown in Fig. 8(d), the model successfully recovers fine image details and markedly mitigates color deviation. Table 5 demonstrates that our method attains state-of-the-art reconstruction metrics.
The effectiveness of dual domain loss. We evaluate the contribution of different domain loss in Table 4. Training RDDM with supervision solely in the RAW domain yields satisfactory performance in image generation metrics but performs poorly in both fidelity and perceptual quality metrics, and also introduces localized color deviation issues, as illustrated in Fig. 9(a). In contrast, training RDDM exclusively in the sRGB domain results in sub-optimal performance across all metrics, although it mitigates the color deviation problem, as shown in Fig. 9(b). Training RDDM with supervision in both the RAW and sRGB domains leads to substantial improvements in fidelity and image perceptual quality metrics. However, certain image generation metrics, such as MUSIQ and CLIPIQA, experience a slight decline, while there is a significant improvement in visual quality, as depicted in Fig. 9(c).
The comparison of text prompt extractors. Ultimately, we undertake a systematic evaluation of alternative text prompt extractors. We devise three distinct text extraction strategies. (1) We directly apply a pre-trained sRGB text prompt extractor (sRGBPE) to the sensor RAW images. (2) ISPPE first converts the sensor RAW data to sRGB via an ISP and then extracts prompts with the same pre-trained sRGB extractor. (3) Our proposed RPE employs DDNet for joint denoising and demosaicking, followed by a PTP module that maps the result into the sRGB domain before prompt extraction. As shown in Table 6, RPE reliably extract text prompts, effectively activating the priors encoded in the pre-trained diffusion network and yielding consistent improvements on image fidelity metrics such as PSRN and image generation metrics such as DISTS, FID, NIQE, MUSIQ and CLIPIQA.
5 Conclusion
We propose RDDM, a novel paradigm for Real-IR, restores directly from the sensor RAW. RDDM exploits the unprocessed, detail-rich signals in sensor RAW data to achieve high fidelity and perceptual quality, thereby alleviating the sub-optimal performance commonly observed in existing Real-IR models that rely on lossy sRGB imagery. Furthermore, RDDM is compatible with diverse Bayer pattern sensor RAW and can be extended to multi-frame input scenarios. The proposed paradigm is applicable to tasks for which the RAW data is accessible, including image and video capture on edge devices. We hope that this work facilitates practical applications and following studies bridging generative modeling, image processing of RAW and image restorations.